find the magnitude and direction of the vector product of two vectors of 10 units each, which are inclined at 30 degrees to each other.
1
Expert's answer
2013-01-10T09:04:20-0500
Vector product.
magnitude = A*B*sin(angle) = 10 * 10 * sin(30 deg) = 10 * 10 * 0.5 = 50;
direction: perpendicular to the plane in which these two vectors lie.